OpenStudy (yaihlyn):

I have a question on Graphing linear inequalities. I dont remember how to do them..

OpenStudy (welshfella):

what's the question?

OpenStudy (yaihlyn):

The graph shows the solution to which system of equations? A) y = x + 2 and y = -x - 4 B) y = 3x + 7 and y = x + 12 C) y = 2x - 3 and y = −2x + 1 D) y = -x + 4 and y = −2x + 5